y=3^(2x+1)
y=9^(4-2x)
solve the following system

Answer by lwsshak3(11628)

y=3^(2x+1)
y=9^(4-2x)
solve the following system
3^(2x+1)=9^(4-2x)
3^(2x+1)=3^2(4-2x)
2x+1=8-4x
6x=7
x=7/6